OpenStudy (anonymous):

Solve the equation for y and then find the value of y for the given x-value . 5x+y=10 x=-2 y= ?

OpenStudy (anonymous):

Okay so if x=-2 you would have to plug it into your equation that is given. 5(-2)+y=10

OpenStudy (jdoe0001):

ahemm... can it be done that way? yes were those the instructions? no

OpenStudy (anonymous):

Then you will have to multiply the numbers together 5(-2)=-10.

OpenStudy (anonymous):

What do you mean? @jdoe0001

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ \textit{Solve the equation for y}}} {\color{red}{ \textit{ and then} }} \)find the value of y for the given x-value .

OpenStudy (anonymous):

Well I was just showing the easier way of doing it but I do know that way

OpenStudy (anonymous):

do u know the answer @Softballstar201

OpenStudy (anonymous):

Okay so 5(-2) = -10 then the equation would be y-10=10

OpenStudy (anonymous):

okay so the answer is -2 @Softballstar201

OpenStudy (anonymous):

Then you would simplify on both sides making it y= 10+10 which would make it y=20

OpenStudy (anonymous):

No it wouldn't be -2 because the 5(-2) is 5x. You gave me x so I plugged in -2 for x

OpenStudy (anonymous):

The answer is y=20 @mamas

OpenStudy (anonymous):

oh thanks can you help me with more ? @Softballstar201

OpenStudy (anonymous):

Sure I have time

OpenStudy (anonymous):

ok give me one sec and ill tag u @Softballstar201

OpenStudy (anonymous):

Alright

DivineSolar (divinesolar):

Could have been easier if you just added 10 to both sides :/